### v1

@@ 4 Proof of Theorem 1
When $\delta$ is above some cutoff, inequality (4.4) is true only if there exists $i\in D(\mathbf{h^{*}})$ such that $i>j$, which contradicts the definition of $j$ under which there exists no $\mathbf{h}\in H_{\sigma}$ such that $h_{i}\geq 1$ for any $i>j$. When $\delta$ is above some cutoff, this contradiction rules out strict equilibria with monotone belief maps where signals other than $s_{1}$ are disclosed.
5 Application: Beliefs in a Just World
We now present an application of our model to the problem of influencing future generations’ beliefs about the returns to effort in the style of Benabou and Tirole (2006). The key difference is that in our model, the strategic choice of the old generation is which verifiable signals to disclose to the young. This is relevant because in our overlapping generations setting, there are potentially many signals that are available for transmission, which can be interpreted as various anecdotes that they have experienced or have heard from their predecessors.
In each period $t\in\mathbb{Z}$, there is a unit mass of young agents and a unit mass of old agents. A representative young agent works in period $t$ by choosing an action $a^{t}\in\mathbb{R}$ given some private signal $s^{t}\in S$ and with probability $\delta\in(0,1)$, the signals disclosed to him by the old agent, i.e., the agent who was young in period t1t-1.999When not in the informative e interpret this as a continuum of dynasties where the old agent in a dynasty passes on information at time $t$ to the young agent in that dynasty. The signals are informative about a constant state of the world $\theta\in\{\underline{\theta},\overline{\theta}\}$, which captures the true returns to effort. We assume that $0<\underline{\theta}<\overline{\theta}<1$. The prior probability that $\theta=\overline{\theta}$ is $\pi_{0}\in\left(0,1\right)$. The old agent chooses which signals from his history to disclose to the young agent. Within each period $t$ the timing is as follows.
  1. 1.
    The old agent begins period $t$ with a history of verifiable signals. He chooses which subset of these signals to disclose to the young agent.
  2. 2.
    The young agent observes the disclosed history with probability $\delta$ and then privately observes a new verifiable signal $s^{t}\in S$ about the state $\theta$.
  3. 3.
    The young agent chooses effort $a^{t}\in\mathbb{R}$.
  4. 4.
    Output $y^{t}\in\{0,1\}$ is realized. Redistribution and consumption occur according to some exogenous redistribution rule.
  5. 5.
    The young agent becomes the old agent in the next period. He has access to his own private signal and with probability $\delta$, the signals passed on to him from the older generation.
As in Piketty (1995) and Benabou and Tirole (2006), a representative young agent in period $t$ produces output according to the following production technology
$y^{t}=\left\{\begin{array}[]{ccl}1&\text{ \ with probability}&\theta a^{t}\\ 0&\text{ \ with probability }&1-\theta a^{t}\end{array}\right$$\text{,}$
where $a^{t}$ is the level of effort chosen by young agent in period $t$.
There is a fixed tax rate denoted by $\tau$ in the economy, and everybody gets a fixed, basic transfer $b$. The period $t$ young agent obtains utility
$u\left(y^{t}\left(1-\tau\right)+b\right)-\frac{1}{2}\left(a^{t}\right)^{2}\text{,}$
where $\frac{1}{2}\left(a^{t}\right)^{2}$ is the cost of effort. [^10] Our conclusions in this section extend qualitatively to more general convex effort cost functions. Each old agent obtains utility $u\left(b\right)+\lambda v(y^{t})$, where $\lambda>0$ captures the old agents’ altruism motives in the sense that they benefit from the consumption of their children. We assume that $u(\cdot)$ is strictly increasing and concave, so that agents are risk-averse. Each agent’s lifetime utility equals the sum of his utilities across the two periods.
A young agent’s belief about $\theta$, before observing his private signal, but after seeing the history $\mathbf{h}\in H$ disclosed by the old agent, is denoted by $\pi(\mathbf{h})$. Denote by $\pi(\mathbf{h},s_{i})$ the young agent’s posterior belief about $\theta$ after he also observes private signal $s_{i}$, so that
$\pi\left(\mathbf{h},s_{i}\right)=\frac{\pi\left(\mathbf{h}\right)\overline{f_{i}}}{\pi\left(\mathbf{h}\right)\overline{f_{i}}+\left(1-\pi\left(\mathbf{h}\right)\right)\underline{f_{i}}}\text{.}$
Let $\widehat{\theta}\left(\pi\right)=\pi\overline{\theta}+\left(1-\pi\right)\underline{\theta}$ denote the expected value of the state conditional on an agent having posterior belief $\pi$. We let $\widehat{\theta}\left(\mathbf{h},s\right)=\widehat{\theta}\left(\pi\left(\mathbf{h},s\right)\right)$ denote the agent’s expectation of the state after he observes the disclosed history as well as his private signal.
Fix a communication success rate $\delta\in(0,1)$ and a constant threshold strategy with cutoff $j\in\{0,1,\dots,l\}$, so that the stationary distributions of histories conditional on states $\overline{\theta}$ and $\underline{\theta}$ are given by $\overline{\mu}$ and $\underline{\mu}$, respectively. We later confirm that these strategies indeed form an equilibrium.
Given information $(\mathbf{h},s)$, the young agent in period $t$ chooses $a^{t}$ to maximize expected utility
$\max_{a^{t}\in\mathbb{R}}\widehat{\theta}\left(\mathbf{h},s_{i}\right)a^{t}u\left(1-\tau+b\right)+\left(1-\widehat{\theta}\left(\mathbf{h},s_{i}\right)a^{t}\right)u(b)-\frac{1}{2}\left(a^{t}\right)^{2}\text{.}$
Let the difference in consumption utility between success and failure be $\Delta c(b)=u\left(1-\tau+b\right)-u(b)$. The first-order condition for an interior solution implies the following linear rule in choosing effort
$a^{t}(\mathbf{h},s_{i})=\widehat{\theta}\left(\mathbf{h},s_{i}\right)\Delta c(b)\text{.}$ (5.1)
In a stationary, symmetric equilibrium all young agents take an action that is a function of $(\mathbf{h},s_{i})$. We assume that the law of large numbers holds for the continuum and that the distribution of agent histories and signals in period $t$ matches the stationary steady-state distribution so that the stationary expected income in the economy is
$Y$ $=$ $\mathbb{E}_{\theta,\left(\mathbf{h},s_{i}\right)}\left[\widehat{\theta}(\mathbf{h},s_{i})a_{k}^{t}(\mathbf{h},s_{i})\right]$
$=$ $\mathbb{E}_{\theta,\left(\mathbf{h},s_{i}\right)}\left[\widehat{\theta}(\mathbf{h},s_{i})^{2}\right]\Delta c(b)\text{,}$
where conditional on the states $\overline{\theta}$ and $\underline{\theta}$, the distributions of histories $\mathbf{h}$ are $\overline{\mu}$ and $\underline{\mu}$, respectively. We also assume that government transfers are ex-ante budget balanced so that $\tau Y=2b+\psi x$, where $x$ is the amount spent on public goods and, like the tax rate $\tau$, is exogenous.
The fact that government budgets are ex-ante balanced means that the value of the basic transfer $b$ does not provide any additional information about the state of the world to the agents.
Let $\kappa\!=\mathbb{E}_{\theta,\left(\mathbf{h},s_{i}\right)}\left[\widehat{\theta}(\mathbf{h},s_{i})^{2}\right]$ and observe that since $b=\left(\tau Y-\psi x\right)/2$, the stationary expected income $Y$ satisfies the following equation in equilibrium
$Y$ $=$ $\kappa\Delta c\left(\frac{\tau Y-\psi x}{2}\right).$ (5.2)
Because $u$ is strictly concave, the right-hand side is decreasing in $Y$ while the left-hand side is strictly increasing. Hence, there is at most one fixed point. Furthermore, since at $Y=0$, the left-hand side is smaller, while at $Y=\max\left\{1,\kappa u(1-\tau)\right\}$, the left-hand side is bigger than the right-hand side. Thus, the intermediate value theorem implies the existence of a fixed point for $Y$ that satisfies (5.2).
At this value of $Y$, we have preferences for the agents that satisfy the assumptions of our model from Section 2. Observe that period $t$ effort by the young agents is increasing in their belief that the state $\theta=\overline{\theta}$, and that in turn the old generation’s welfare is increasing in this effort since $Y$ is increasing in $a^{t}$. Note that since agents are atomistic, if $\lambda=0$ an old agent is not directly impacted by the actions of the young agent from the same dynasty. If $\lambda>0$, the old agent in period $t$ can be thought of as caring about the young agent’s utility but discounting the effort cost, as in Benabou and Tirole (2006), or just enjoying a fraction of the young person’s income. These preferences motivate the old agent to convince the young that returns to effort are high, which is what we assumed in Section 2.
Proposition 1 and Theorem 2 then imply that an equilibrium in constant threshold strategies exists for all $\delta$. We let $j$ denote a threshold that can arise in one of those equilibria and note that at this value of $\delta$ and $j$, the above effort choices are consistent with the average income $Y$ that balances ex ante the government budget via the above fixed-point argument.
Note that this setup shares a key feature of Piketty (1995) and Benabou and Tirole (2006), namely that holding fixed the fundamentals (such as the signal structure $\mathbf{f}$, agents’ preferences, and prior belief $\pi_{0}$), there are multiple steady-state “beliefs”, which are due to different possible equilibria (i.e., intergenerational disclosure conventions. In particular, there is always an uninformative steady-state equilibrium in which old agents disclose no signals (i.e., the threshold is $j=0$). In this equilibrium, young agents’ public belief remains at the prior and their effort responds only to the privately observed signal. Proposition 1 implies that for any $\delta$ there also exists a steady-state equilibrium in which the only signals ever transmitted are the most favorable ones (i.e., the threshold is $j=1$). For lower $\delta$, Theorem 2 shows that additional equilibria with higher cutoffs ($j>1$). These equilibria correspond to the Benabou and Tirole (2006) discussion of the United States versus Europe. That is, a society in which the norm is that elders pass on only the strongest evidence supporting high returns to effort has more dispersed equilibrium beliefs and higher equilibrium effort (a belief in the “American Dream”), than a society with more muted disclosure norms.
Comparative Statics on $\delta$:
Consider a planner who cares about maximizing the sum of utilities of the young and old agent. When writing the planner’s problem, we ignore the preferences for the old over the income of the young by considering the case where $\lambda=0$. Old agents in society would still want to convince the young that returns to effort are high, in order to motivate effort and have a higher basic transfer, if there was a finite number of dynasties.
Like in Benabou and Tirole (2006), there is a natural rate of information recall, $\delta_{0}$, but this could be increased at a cost $x$. Earlier in Section 5, we saw that $x$ was the amount spent on provision of public goods and we interpret it now as the amount spent on education. Spending more on education increases the effective $\delta$ in society. We assume that $\delta\left(x\right)=1-\left(1-\delta_{0}\right)e^{-x}$.
The planner’s problem is to choose $\delta$ to maximize
$\max_{\delta}\;\mathbb{E}_{\theta,(\mathbf{h},s_{i})}\Big[u\big($ $(1-\tau)\,y(\mathbf{h},s_{i})+b\big)-\tfrac{1}{2}\,a(\mathbf{h},s_{i})^{2}+u(b)\Big]$
$\text{s.t.}\quad a(\mathbf{h},s_{i})$ $=\widehat{\theta}(\mathbf{h},s_{i})\,\Delta c(b),$
$y(\mathbf{h},s_{i})$ $=\theta\,a(\mathbf{h},s_{i}),$
$b$ $=\frac{\tau Y-\psi x}{2},\qquad\text{where}\quad Y=\mathbb{E}_{\theta,(\mathbf{h},s_{i})}\!\left[\widehat{\theta}(\mathbf{h},s_{i})^{2}\right]\Delta c(b).$
With $\psi=0$, the planner is free to adjust $\delta$ in any way she likes. Because expected output $Y$ is increasing in $\kappa$ and thus the variance in the agents posterior over $\theta$, Proposition 2 implies that the planner always prefers a strictly larger $\delta$. For sufficiently high $\psi$, the optimal choice of the planner is to leave $\delta$ at $\delta_{0}$. While for intermediate cases, some intermediate $\delta$ will be optimal.
6 Concluding Remarks

### v2

@@ 4 Proof of Theorem 1
Recall our hypothesis that $j\geq 2$, which by our earlier conclusion, implies that $D(\mathbf{h^{*}})$ is non-empty and does not contain anything strictly greater than $j$. Since $S$ is a finite set, we know that there exists $\delta^{*}\in(0,1)$ such that for every $\delta>\delta^{*}$, inequality (4.2) fails for any $j$ and non-empty $D(\mathbf{h^{*}})$ that satisfy $j\geq 2$ and $j\geq\max D(\mathbf{h^{*}})$. When $\delta>\delta^{*}$, the above contradiction rules out strict monotone equilibria where signals other than $s_{1}$ are disclosed on the equilibrium path.
Strict Threshold Equilibria:
Suppose for some $\delta$ is close to $1$, there exists a strict threshold equilibrium where agents use a threshold strategy $\sigma$ under which $j\geq 2$. We establish the following lemma, which is the analog of Lemma 2 when the solution concept is threshold equilibrium.
  Lemma 3.
Fix any threshold equilibrium. If an agent’s history is $\mathbf{h^{*}}$, then his immediate predecessor’s history is either $\mathbf{h^{*}}$ or $\mathbf{h^{*}}-\mathbf{1}_{j}$.
The proof is in Appendix A.5.
$\mu_{\theta}(\mathbf{h^{*}})$ $=\mu_{\theta}(\mathbf{h^{*}}-\mathbf{1}_{j})\cdot\delta f_{\theta,j}\cdot\sum_{k=0}^{\infty}\delta^{k}\left(1-\sum_{i=1}^{j}f_{\theta,i}\right)^{k}$
$=\mu_{\theta}(\mathbf{h^{*}}-\mathbf{1}_{j})\cdot\frac{\delta f_{\theta,j}}{1-\delta\left(1-\sum_{i=1}^{j}f_{\theta,i}\right)}.$ (4.3)
After receiving private signal $s_{j}$ at history $\mathbf{h^{*}}-\mathbf{1}_{j}$, an agent has an incentive to disclose history $\mathbf{h^{*}}$ to his successor only if $\pi(\mathbf{h^{*}})\geq\pi(\mathbf{h^{*}}-\mathbf{1}_{j})$, which according to (4), translates into
When $\delta$ is above some cutoff, inequality (4.4) fails for any $j\geq 2$, which leads to a contradiction.
5 Discussion
In Section 5.1, we discuss the connections between our overlapping generations model and one-shot disclosure games between a sender and a receiver where the number of signals the sender has follows a geometric distribution. In Sections 5.2 and 5.3, we study two extensions of our baseline model, one where agents face a constraint on the maximal number of signals they can disclose and one where the communication success rate depends on the set of signals transmitted.
5.1 Comparison to one-shot disclosure games
A natural question is whether our conclusions extend to a one-shot disclosure game between a sender and a receiver where the sender has an unknown number of signals. Suppose the state is binary $\theta\in\{\overline{\theta},\underline{\theta}\}$. The sender wants to maximize the probability the receiver’s belief assigns to state $\overline{\theta}$. The number of verifiable (conditionally independent) signals the sender has is drawn from a geometric distribution with parameter $\delta$. The receiver only observes the number of each signal the sender discloses, so that each receiver’s information set belongs to the set $H\equiv\mathbb{N}^{l}$.
In this game, the sender’s strategy is $\sigma_{D}:H\rightarrow H$ subject to the constraint that $\sigma_{D}(\mathbf{h})\leq\mathbf{h}$ for every $\mathbf{h}\in H$. For any $\mathbf{h}\in H$ and $i\in\{0,1,...,l\}$, recall that $\mathbf{h}[i]\in H$ is defined as a vector where the first $i$ entries coincide with $\mathbf{h}$ and the other entries are zero. We say that $\sigma_{D}$ is a threshold strategy if there exists $J:H\rightarrow\{0,1,...,l\}$ such that $\sigma_{D}(\mathbf{h})=\mathbf{h}[J(\mathbf{h})]$ for every $\mathbf{h}\in H$. A threshold strategy is a constant threshold strategy if $J(\mathbf{h})=J(\mathbf{h^{\prime}})$ for every $\mathbf{h}\neq\mathbf{h^{\prime}}$. The receiver’s belief map is $\pi:H\rightarrow[0,1]$. The monotonicity property of the belief map is defined in the same way as in our overlapping generations model. The solution concept is Perfect Bayesian equilibrium. Our definitions of monotone equilibrium, threshold equilibrium, and constant threshold equilibrium extend naturally to this one-shot game.
Statement 1 of our Theorem 1, which concerns strict monotone equilibria, fails to extend to this one-shot game, thereby distinguishing the one-shot game from our overlapping-generations game. Intuitively, this is because the gradualism property stated in Section 4 breaks down, which means that Lemma 2 is no longer true. In the one-shot game, the sender will observe every combination of signals with positive probability, as opposed to only a subset of those combinations determined by players’ equilibrium strategy in our overlapping generations game.
In what follows, we construct strict monotone equilibria of the one-shot game under which signals other than $s_{1}$ is disclosed with positive probability. We introduce a class of strategies for the sender parameterized by integer $k\in\mathbb{N}$.
  • Strategy $k$: For every $j\neq k+1$, if the sender observes $j$ copies of signal $s_{1}$, then he discloses all the $s_{1}$ he has and nothing else. If the sender has $k+1$ signal $s_{1}$ but no $s_{2}$, then he discloses $k$ signal $s_{1}$ and nothing else. If the sender has $k+1$ signal $s_{1}$ and at least one $s_{2}$, then he discloses $k+1$ signal $s_{1}$, one $s_{2}$, and nothing else.
Strategy $k$ is not a threshold strategy. This is because when the sender has $k+1$ signal $s_{1}$ and two $s_{2}$, he discloses only one $s_{2}$ and conceals the other. If the sender uses Strategy $k$ in the one-shot game, $s_{2}$ will be disclosed at some positive-probability information set since the sender will observe $k+1$ signal $s_{1}$ and one $s_{2}$ with positive probability. However, if agents use Strategy $k$ in our overlapping generations game, $s_{2}$ will not be disclosed on the equilibrium path as history $(k+1)\mathbf{1}_{1}$ will occur with zero probability.
  Proposition 3.
For any $\delta\in(0,1)$, there exists $k\in\mathbb{N}$ such that there exists a strict monotone equilibrium in the one-shot disclosure game where the sender uses Strategy $k$.
The proof is in Appendix A.6.
In contrast, Statement 2 of Theorem 1 which concerns strict threshold equilibria, can be extended to this one-shot disclosure game.
  Proposition 4.
There exists $\delta^{*}\in(0,1)$ such that when $\delta>\delta^{*}$, the sender will never disclose $s_{2},...,s_{l}$ in any strict threshold equilibrium.
The proof is in Appendix A.7.
Theorem 2, which focuses on constant-threshold equilibria, also extends to this one-shot disclosure game. This is because when the sender uses a constant disclosure threshold $n$, the only relevant incentive constraint is for him to disclose signal $s_{n}$ whenever it is available, i.e., $\pi(\mathbf{h}[n])\leq\pi(\mathbf{h}[n]+\mathbf{1}_{n})$ for every $\mathbf{h}\in H$. This is the case if and only if inequality (3.8) holds. The rest of argument follows from the proof of Theorem 2 in Section 3, which is omitted.
5.2 Extension to disclosure constraints
We now discuss two extensions of our baseline model where our main result Theorem 1 extends. In our baseline model, agents can disclose any number of available signals. In practice, agents may face constraints on the number of signals they can disclose. Consider an extension where each agent can disclose at most $N\in\mathbb{N}$ signals. That is, after receiving private signal $s_{i}$ at history $\mathbf{h}$, he can disclose $\mathbf{h^{\prime}}=(h_{1}^{\prime},...,h_{l}^{\prime})$ if and only if $\mathbf{h^{\prime}}\leq\mathbf{h}+\mathbf{1}_{i}$ and $\sum_{j=1}^{l}h_{j}^{\prime}\leq N$.
Our proof of Theorem 1 extends to any value of $N$. To see this, notice that the gradualism property, the improvement principle, and the no-turning-back property remain true regardless of $N$. Lemma 2 and Lemma 3 also extend since their proofs only need to rule out the possibility that the current-period history is $\mathbf{h^{*}}$ yet the previous-period history is neither $\mathbf{h^{*}}$ nor $\mathbf{h^{*}}-\mathbf{1}_{j}$. Such a possibility is also ruled out when the agents face an additional constraint when they decide which signals to disclose. The calculations after Lemma 2 and Lemma 3 concern the probability of history $\mathbf{h^{*}}$, which by construction, must satisfy our additional constraint. Equations (4.1) and (4) express the probability of $\mathbf{h^{*}}$ as a function of the probability of history $\mathbf{h^{*}}-\mathbf{1}_{j}$. Despite the disclosure set $D(\mathbf{h^{*}})$ may depend on $N$, the way we derived a contradiction only uses a particular property of $D(\mathbf{h^{*}})$, which is $j\geq\max D(\mathbf{h^{*}})$, which holds regardless of $N$.
Nevertheless, Theorem 1 characterizes a common property of all strict monotone equilibria and strict threshold equilibria when $\delta$ is large. The set of strict monotone equilibria and the set of strict threshold equilibria will depend on the upper bound $N$. For example, disclosing all realized $s_{1}$ will not be an equilibrium since the agents’ strategy will violate our new constraint.
5.3 Extension to message-dependent communication success rate
In our baseline model, the probability that communication succeeds is independent of the set of signals an agent chooses to disclose to their successor. In practice, however, the success rate may depend on the number of the signals disclosed. For example, children may absorb information from their parents with lower probability when parents attempt to convey too much.
Suppose now that when agent $t$ discloses $\mathbf{h}$, the probability that agent $t+1$ receives $\mathbf{h}$ is $\delta(\mathbf{h})\equiv\exp(-\Delta\cdot p(\mathbf{h}))$ where $\Delta>0$ is a parameter and $p:H\rightarrow[\underline{p},\overline{p}]$ for some $0<\underline{p}<\overline{p}<\infty$ such that $p(\mathbf{h})$ depends on $\mathbf{h}$ only through its length and $p(\mathbf{h})$ weakly increases as the length of $\mathbf{h}$ increases. In this setting, communication is more likely to fail when the history disclosed is longer. As $\Delta\rightarrow 0$, communication success rate converges to $1$.
Theorem 1 extends to this more general setting. To see this, notice that the gradualism property and the no-turning-back property hold independently of the communication success rate. A modified version of the improvement principle holds in this case, that when $\mathbf{h}^{t+1}$ is disclosed when the current-period history is $\mathbf{h}^{t}$, the expected value of $v(\theta,a^{*}(\pi))$ when the agent discloses $\mathbf{h}^{t+1}$ is greater than the expected value when he discloses $\mathbf{h}^{t}$. We explain why Lemma 2 extends by the end of Appendix A.4. The same logic applies to the proof of Lemma 3, which we omit in order to avoid repetition.
In order to show that agents will never disclose any $s_{2},...,s_{l}$ in any strict monotone equilibrium when $\Delta$ is small enough, we need to modify equation (4.1) as
$\mu_{\theta}(\mathbf{h^{*}})=\mu_{\theta}(\mathbf{h^{*}}-\mathbf{1}_{j})\cdot\delta(\mathbf{h^{*}})\cdot f_{\theta,j}\cdot\sum_{n=0}^{\infty}\delta(\mathbf{h^{*}})^{n}\cdot(1-\sum_{i\in D(\mathbf{h^{*}})}f_{\theta,i})^{n}.$ (5.1)
This is because when the current-period history is $\mathbf{h^{*}}-\mathbf{1}_{j}$, the history in the next period is $\mathbf{h^{*}}$ if and only if the current-period private signal is $s_{j}$, the agent discloses $\mathbf{h^{*}}$, and communication succeeds. Conditional on state $\theta$, the above event occurs with probability $\delta(\mathbf{h^{*}})\cdot f_{\theta,j}$. Moreover, when the current history is $\mathbf{h^{*}}$, the next period’s history remains at $\mathbf{h^{*}}$ if and only if (i) the agent’s private signal does not belong to $D(\mathbf{h^{*}})$ in which case he discloses $\mathbf{h^{*}}$ to the next agent and (ii) communication succeeds. Conditional on state $\theta$, the above event occurs with probability $\delta(\mathbf{h^{*}})\cdot(1-\sum_{i\in D(\mathbf{h^{*}})}f_{\theta,i})$. Equation (5.1) implies that $\pi(\mathbf{h^{*}})<\pi(\mathbf{h^{*}}-\mathbf{1}_{j})$ if and only if
$L_{j}\cdot\frac{1-\delta(\mathbf{h^{*}})\cdot(1-\sum_{i\in D(\mathbf{h^{*}})}\underline{f_{i}})}{1-\delta(\mathbf{h^{*}})\cdot(1-\sum_{i\in D(\mathbf{h^{*}})}\overline{f_{i}})}<1.$ (5.2)
Since $\delta(\mathbf{h^{*}})=\exp(-\Delta\cdot p(\mathbf{h^{*}}))$ and the function $p(\cdot)$ is uniformly bounded from above and from below, we know that there exists $\Delta^{*}>0$ such that when $\Delta<\Delta^{*}$, inequality (5.2) holds for all $j$ and $D(\mathbf{h^{*}})$ that satisfy $j\geq 2$ and $j\geq\max D(\mathbf{h^{*}})$. After receiving private signal $s_{j}$ at history $\mathbf{h^{*}}-\mathbf{1}_{j}$, (i) by disclosing $\mathbf{h^{*}}$, an agent induces public belief $\pi(\mathbf{h^{*}})$ with probability $\delta(\mathbf{h^{*}})$ and induces public belief $\pi(\mathbf{0})$ with probability $1-\delta(\mathbf{h^{*}})$; (ii) by disclosing $\mathbf{h^{*}}-\mathbf{1}_{j}$, an agent induces public belief $\pi(\mathbf{h^{*}}-\mathbf{1}_{j})$ with probability $\delta(\mathbf{h^{*}}-\mathbf{1}_{j})$ and induces public belief $\pi(\mathbf{0})$ with probability $1-\delta(\mathbf{h^{*}}-\mathbf{1}_{j})$. Disclosing $\mathbf{h^{*}}-\mathbf{1}_{j}$ leads to a strictly higher expected payoff than disclosing $\mathbf{h^{*}}$ when $\Delta$ is small enough since $\pi(\mathbf{h^{*}})<\pi(\mathbf{h^{*}}-\mathbf{1}_{j})$, $\delta(\mathbf{h^{*}})\leq\delta(\mathbf{h^{*}}-\mathbf{1}_{j})$, and from the improvement principle $\pi(\mathbf{0})\leq\min\{\pi(\mathbf{h^{*}}),\pi(\mathbf{h^{*}}-\mathbf{1}_{j})\}$. This leads to a contradiction, which rules out strict monotone equilibria where agents disclose signals other than $s_{1}$. The same reasoning can be applied to show that agents will never disclose $s_{2},...,s_{l}$ in any strict threshold equilibrium when $\Delta$ is small enough, by replacing $\sum_{i\in D(\mathbf{h^{*}})}f_{\theta,i}$ in (5.1) and (5.2) with $\sum_{i=1}^{j}f_{\theta,i}$.
6 Conclusion